# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v1

@@ Abstract
In this paper, we propose Text-based Open Molecule Generation Benchmark (TOMG-Bench), the first benchmark to evaluate the open-domain molecule generation capability of LLMs. TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each task further contains three subtasks, with each subtask comprising 5,000 test samples. Given the inherent complexity of open molecule generation, we have also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules. Our comprehensive benchmarking of 25 LLMs reveals the current limitations and potential areas for improvement in text-guided molecule discovery. Furthermore, with the assistance of OpenMolIns, a specialized instruction tuning dataset proposed for solving challenges raised by TOMG-Bench, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench. Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
@@ 1 Introduction
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals Keiser et al. (2010) to materials science Higuchi et al. (2023). Normally, molecule discovery is a trial and error process Ekins (2024), which requires repetitive experimentation and data analysis Mattern and Grosser (2023). Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market Lee et al. (2018).
With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) Wang et al. (2023), there has been a significant step forward. As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions. With the assistance of GNNs, researchers could analyse the properties of molecules Cai et al. (2022) and generate new molecule candidates Jin et al. (2018). However, challenges still exist. GNN-based methods struggle to generalize to different tasks Chen et al. (2024), necessitating costly data collection and preparation for different downstream tasks. Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design Li et al. (2024c).
@@
In Contrast, Large Language Models (LLMs) have shown their great generalization capability Achiam et al. (2023) and could be easily adapted to different research fields. For instance, Cascella et al. (2023) utilizes ChatGPT for supporting clinical practice and Zhen et al. (2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
@@
As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound Weininger (1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages. With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain Li et al. (2024b). The generalization capability makes LLMs a viable option in molecule discovery. Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists Zhang et al. (2024); Li et al. (2024d). They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
@@
While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging Li et al. (2024a). A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation. Although the task of molecule-caption translation Edwards et al. (2022) is crucial for bridging the gap between the molecular and textual domains, it still has several limitations that need to be addressed:
@@
On the one hand, there is a concern about the generalization of the molecule-caption translation task Li et al. (2024b). In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task. In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples Li et al. (2024b). This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well. On the other hand, a critical issue is the inability to propose new molecule structures. The ultimate goal of molecule discovery is not just to understand and describe existing chemical compounds but to innovate and discover new ones, particularly in the context of drug discovery, indicating that the current molecule-caption translation task and the corresponding evaluation metrics fall short in this regard. Therefore, addressing these challenges is essential for harnessing the full potential of LLMs in molecule discovery.
@@
In our efforts to bridge the gap between the natural language and the molecular spaces and to further facilitate LLMs as chemist assistants in molecule discovery, we propose a novel benchmark, Text-based Open Molecule Generation Benchmark (TOMG-Bench). TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations. As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exact matched molecule. Instead, we adopt chemical toolboxes like RDKit Landrum (2013) to test whether the generation meets the requirements. In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them. Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive and robust assessment of whether LLMs truly grasp the molecular space. Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules. Moreover, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules from an existing molecule database. OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and extra-large) to tailor different training purposes.
To encapsulate our contributions, they are primarily threefold:
  • Firstly, we are the first to propose the task of Text-based Open Molecule Generation, along with a new set of evaluation metrics to benchmark the performance of LLMs in this task.
  • Secondly, we have benchmarked a suite of 25 LLMs, shedding light on the limitations of the targeted molecule generation.
@@
  • Thirdly, we introduce an instruction tuning set, OpenMolIns, which helps Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
@@ 2 Related Work
In this section, we briefly review related work about developing Artificial Intelligences (AIs) in Molecule Discovery and, more specifically, in text-based molecule generation tasks.
@@ 2.1 Development of AIs in Molecule Discovery
Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design Du et al. (2022). Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules. Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods Wigh et al. (2022); Wu et al. (2018); Zhou et al. (2023). For instance, MolReGPT  Li et al. (2024b) leverages large language models (LLMs) like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks. Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) Peng and Zhao (2019); Le et al. (2019), Recurrent Neural Networks (RNNs) Grisoni et al. (2020); Popova et al. (2019), Graph Neural Networks (GNNs) Wang et al. (2023); Sun et al. (2022), and Transformer-based networks Xia et al. (2023); Balaji et al. (2023); Edwards et al. (2022).
@@ 2.2 Text-based Molecule Generation
Text-based Molecule Generation (Text2Mol) Edwards et al. (2021) has recently emerged as a transformative approach to molecule discovery. This task centres on retrieving molecules using natural language descriptions as search queries, requiring the creation of paired datasets of molecules and their corresponding textual representations. This enables the learning of a shared semantic embedding space for efficient retrieval. Early approaches leveraged transformer-based models like MolT5 Edwards et al. (2022), employing self-supervised learning on large datasets to generate high-quality Simplified Molecular Input Line Entry System (SMILES) strings from textual inputs. Subsequent advancements, such as KV-PLM Zeng et al. (2022), MoMu Su et al. (2022), and BioT5 Pei et al. (2023), integrated molecular graphs and biochemical text to improve both understanding and generation capabilities. 3D-MoLM Li et al. (2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations. The application of large language models (LLMs) like MolReGPT Li et al. (2024b) and ICMA Li et al. (2024a) as in-context learners has also shown significant promise. These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context. Most recently, MolReFlect Li et al. (2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively. Unlike this targeted generation task, in this paper, we propose a Text-based Open Molecule Generation task to enable LLMs to generate an exactly matched molecule rather than set a specific target.
@@ 3 TOMG-Bench
In this section, we propose TOMG-Bench to comprehensively assess the performance of LLMs in molecular space. Specifically, the benchmark is composed of three basic tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks. Additionally, we have created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
Figure 2: Data construction workflow and evaluation process of TOMG-Bench.
@@ 3.1 Dataset Categorization
The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want. Following the difficulty of the tasks, we demonstrate the content of the three basic tasks as well as their corresponding subtasks:
@@
MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target. To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics). These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
@@ 3.2 Dataset Construction
Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations. For instance, ChEBI-20  Edwards et al. (2022), a dataset for the molecule-caption translation task, contains only 33,000 samples for training, whereas image captioning datasets like MS COCO  Chen et al. (2015) have over 1,500,000 annotated captions on more than 330,000 images. Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming. In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction. Instead, we focus on basic molecule structural properties and basic molecule operations that could be validated by chemical toolboxes to construct our dataset.
For MolCustom, we randomly generate 5,000 prompts as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups. For MolEdit and MolOpt, we sample molecules from a specific molecule database. Specifically, we select two molecule databases for this work: Zinc-250K Sterling and Irwin (2015), and PubChem  Kim et al. (2019). Zinc-250K has 250,000 molecules, which is smaller than PubChem, which has 10 million molecules. To facilitate the fast calculation of metrics mentioned in Section 3.5, we choose Zinc-250K for sampling the test molecules in TOMG-Bench. Each subtask is allocated 5,000 test samples. After sampling, we utilize RDKit Landrum (2013), a molecular informatics toolbox, to collect basic molecule statistics. There are functions available to calculate the required characteristics, especially the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts. Further details will be provided in the Appendix A.

### v2

@@ Abstract
In this paper, we propose Text-based Open Molecule Generation Benchmark (TOMG-Bench), the first benchmark to evaluate the open-domain molecule generation capability of LLMs. TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each major task further contains three subtasks, while each subtask comprises 5,000 test samples. Given the inherent complexity of open molecule generation evaluation, we also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules. Our comprehensive benchmarking of 25 LLMs reveals the current limitations as well as potential areas for improvement in text-guided molecule discovery. Furthermore, we propose OpenMolIns, a specialized instruction tuning dataset established for solving challenges raised by TOMG-Bench. Fine-tuned on OpenMolIns, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench. Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
@@ 1 Introduction
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023). Normally, molecule discovery is a trial-and-error process (Ekins, 2024), which requires repetitive experimentation and data analysis (Mattern & Grosser, 2023). Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market (Lee et al., 2018).
With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) (Wang et al., 2023), there has been a significant step forward. As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions. With the assistance of GNNs, researchers could analyse the properties of molecules (Cai et al., 2022) and generate new molecule candidates (Jin et al., 2018). However, challenges still exist. GNN-based methods struggle to generalize to different tasks (Chen et al., 2024), necessitating costly data collection and preparation for different downstream tasks. Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design (Li et al., 2024c).
In contrast, Large Language Models (LLMs) have shown their great generalization capability (OpenAI, 2023) and could be easily adapted to different research fields. For instance, Cascella et al. (2023) utilize ChatGPT for supporting clinical practice and Zhen et al. (2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
@@
As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound (Weininger, 1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages. With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain (Li et al., 2024b). The generalization capability makes LLMs a viable option in molecule discovery. Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists (Zhang et al., 2024; Li et al., 2024d). They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
@@
While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging (Li et al., 2024a). A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation. Although the molecule-caption translation datasets including ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023) and L+M-24 (Edwards et al., 2024) are crucial for bridging the gap between the molecular and textual domains, they still face several limitations that need to be addressed:
@@
On the one hand, there is a concern about the generalization of models trained on the molecule-caption translation task (Li et al., 2024b). In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task. In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples (Li et al., 2024b). This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well. On the other hand, a critical issue is the inability to propose new molecule structures. The ultimate goal of molecule discovery is not just to understand and describe existing chemical compounds but to innovate and discover new ones, particularly in the context of drug discovery, indicating that the current molecule-caption translation task and the corresponding evaluation metrics fall short in this regard. Therefore, addressing these challenges is essential for harnessing the full potential of LLMs in molecule discovery.
@@
In our efforts to bridge the gap between the natural language and the molecular spaces and to further facilitate LLMs as chemist assistants in molecule discovery, we propose a novel benchmark, Text-based Open Molecule Generation Benchmark (TOMG-Bench). TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations. As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exactly matched molecule. Instead, we design automatic evaluation processes with chemical toolboxes like RDKit (Landrum, 2013) to test whether the generation meets the requirements. In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them. Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive assessment of whether LLMs truly grasp the molecular space. Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules. To solve the challenges raised by TOMG-Bench, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules with meticulously designed rules from an existing molecule database. OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and xlarge) to study the data scaling law of LLMs in the text-based open molecule generation task.
Our contributions are primarily threefold:
  • 1.
    TOMG-Bench introduces the task of Text-based Open Molecule Generation, which consists of three major tasks and their nine subtasks in total to synthetically benchmark the performance of LLMs in this task, demonstrating the limitations of the targeted molecule generation datasets like ChEBI-20.
  • 2.
    We benchmarked a suite of 25 LLMs, providing a brand-new perspective for the evaluation of LLMs.
  • 3.
@@
    We introduce OpenMolIns, an instruction tuning dataset, enabling Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
@@ 2.1 Development of AIs in Molecule Discovery
Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design (Du et al., 2022). Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules. Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods (Wigh et al., 2022; Wu et al., 2018; Zhou et al., 2023). For instance, MolReGPT  (Li et al., 2024b) leverages LLMs like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks. Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) (Peng & Zhao, 2019; Le et al., 2019), Recurrent Neural Networks (RNNs) (Grisoni et al., 2020; Popova et al., 2019), Graph Neural Networks (GNNs) (Wang et al., 2023; Sun et al., 2022), and Transformer-based networks (Xia et al., 2023; Balaji et al., 2023; Edwards et al., 2022).
@@ 2.2 Text-based Molecule Generation
Text-based Molecule Generation (Edwards et al., 2021) has recently emerged as a transformative task to molecule discovery. Previously, this task centres on generating targeted molecules based on natural language descriptions, requiring the creation of paired datasets of molecules and their corresponding textual representations. Early approaches leveraged transformer-based models like MolT5 (Edwards et al., 2022), employing self-supervised learning on large datasets to generate high-quality SMILES strings from textual inputs. Subsequent advancements, such as KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023), integrated molecular graphs or biochemical text to improve both understanding and generation capabilities. 3D-MoLM (Li et al., 2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations. Meanwhile, the application of LLMs like MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) as in-context learners has also shown significant promise. These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context. Most recently, MolReFlect (Li et al., 2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively.
@@ 3 TOMG-Bench
In this section, we propose TOMG-Bench to comprehensively assess the performance of LLMs in molecular space. Specifically, the benchmark is composed of three basic tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks. Additionally, we created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
Figure 2: Task illustrations of TOMG-Bench. Both molecules to the right of the arrow are acceptable solutions.
@@ 3.1 Dataset Categorization
The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want. As shown in Figure 2, following the difficulty of the tasks, we demonstrate the three basic tasks as well as their corresponding subtasks:
@@
MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target. To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery and can be effectively validated: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics). These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
@@
Compared to previous text-based target molecule generation datasets, the task design in TOMG-Bench is rather naive, but focuses more on the details of the molecular substructure, which puts even more rigorous requirements on both the molecular understanding and generation capabilities of LLMs. Furthermore, the accurate understanding on molecule substructures and editing operations also ensures the explainability and trustworthiness of generated molecules.
@@ 3.2 Dataset Construction
Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations. For instance, ChEBI-20  (Edwards et al., 2022), a dataset for the molecule-caption translation task, contains only 26,407 samples for training. In contrast, datasets for the text-to-image generation task in the computer vision field like MS COCO  (Chen et al., 2015) have over 1,500,000 annotated captions on more than 330,000 images. Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming. In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction. Instead, we focus on basic molecule structural properties and molecule operations that could be validated by chemical toolboxes, which makes it easy to increase the volume of data, thereby resolving data hunger in chemical corpora.
In TOMG-Bench, three major tasks consist of nine subtasks in total. Each subtask is allocated 5,000 test samples. Figure 4 illustrates the data construction process.
For MolEdit and MolOpt, we sample molecules from a specific molecule database. We mainly consider two molecule databases: Zinc-250K (Sterling & Irwin, 2015), and PubChem  (Kim et al., 2019). Zinc-250K has 250,000 molecules, while PubChem has about 10 million molecules. Therefore, the data size of Zinc-250K is more suitable for us to build the test sets, while PubChem could be further sampled to generate the instruction tuning dataset. After sampling, we utilize RDKit (Landrum, 2013), a molecular informatics toolbox, to collect basic molecule statistics, including the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts.
For the MolCustom task, we randomly generate 5,000 instructions as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups. Further details are provided in the Appendix D.